Derive an expression H(ω) in the form of H(ω) = (A + j B) / (C + j D), wherein A, B, C and D are functions of R1, R2, L, and C.

R1 + R2 D(1) C Dout(0) L
Answer #1

Vors2 f M R R2 Į Vo(5) SC Virs) SL Applying nodal analysis at node 0- Vo(s) Vo(s)- Vics) + + = 6 Vors) YSC R R2+ SL - + sct !HIW) = R2+ J WL Rit jw R, R2 C + (w)? RLC that jul How) = R2 + JWL CA R; - w PLC +R2) + jw CRIR2C+L) so A = R2 B = WL c=R,_W?